# Network-Level Prompt and Trait Leakage in Local Research Agents

### v2

@@ Abstract
We show that Web and Research Agents (WRAs)—language model-based systems that investigate complex topics on the Internet—are vulnerable to inference attacks by passive network adversaries such as ISPs. These agents could be deployed locally by organizations and individuals for privacy, legal, or financial purposes. Unlike sporadic web browsing by humans, WRAs visit domains with distinguishable timing correlations, enabling unique fingerprinting attacks.
Specifically, we demonstrate a novel prompt and user trait leakage attack against WRAs that only leverages their network-level metadata (i.e., visited IP addresses and their timings). We start by building a new dataset of WRA traces based on user search queries and queries generated by synthetic personas. We define a behavioral metric (called OBELS) to comprehensively assess similarity between original and inferred prompts, showing that our attack recovers over 73% of the functional and domain knowledge of user prompts. Extending to a multi-session setting, we recover up to 19 of 32 latent traits with high accuracy. Our attack remains effective under partial observability and noisy conditions. Finally, we discuss mitigation strategies that constrain domain diversity or obfuscate traces, showing negligible utility impact while reducing attack effectiveness by an average of 29%.
@@ 1 Introduction
Web and Research Agents (WRAs) are reshaping how users interact with information. Unlike traditional assistants that answer isolated queries, these Large Language Model (LLM)-powered systems autonomously plan, browse, and synthesize knowledge across the web [30]. Major AI providers are integrating such agents directly into their flagship assistants or browsers, including OpenAI [55], Google Gemini [23], Grok [85], Mistral [41], and Perplexity [58]. Open-source counterparts such as LangChain’s Local Deep Researcher [35], GPT Researcher [21], and Agent Laboratory [65] are also widely adopted in academic and developer communities. Their growing deployment in assistants, productivity platforms, and enterprise environments, particularly within privacy-sensitive sectors such as healthcare, law, and finance, raises urgent concerns about their security and privacy implications. While on-device execution through locally-deployable models and TLS-encrypted access could suggest privacy guarantees, this perception is misleading. Yet WRAs leave unavoidable, structured metadata traces that can reveal sensitive user intent, creating an underexplored leakage vector.
As autonomous agent systems, WRAs inherit vulnerabilities from both LLMs and traditional web browsing, such as system prompt extraction [31], model memorization [12], metadata exposure [69], and behavioral fingerprinting [8]. Yet they also expose a distinct and underexplored vulnerability: their behavioral traces. They reflect the agent’s internal reasoning process and can reveal sensitive or stigmatized user intent [52, 73]. Even if prompts and page contents remain fully encrypted, externally visible metadata, such as domain names, access order, payload size, and timing, can silently expose user intent [52, 73]. Because WRAs autonomously execute multi-step research plans [90], they produce cascades of semantically related domain visits in dense temporal bursts [58]. These structured, high-throughput access patterns [17, 42] are unlike human browsing behavior and make WRAs particularly amenable to profiling and inference attacks [17, 42] while introducing new risks through autonomous, high-throughput interaction with live websites.
@@
We adopt a realistic threat model in which a passive adversary, such as an ISP, enterprise firewall, or local network operator, can observe only domain-level traffic metadata, without access to prompts, page contents, or model internals. Despite this extremely limited visibility, we show that adversaries can reliably recover what users asked and who they are.
@@
Motivating Examples. Consider a user seeking reproductive health information: their research agent may visit clinics, support forums, and medical sites. Even without access to page contents, these traces alone can expose the agent’s internal reasoning process and deeply stigmatized intent. As WRAs become integrated into daily decision-making, their behavioral traces present a new leakage channel. We show that WRA visits 70-140 distinct domains for each query in a short period of time, making it an easy target to isolate. In essence, privacy failures emerge not from what agents say, but from how they act.
@@
Our Work. We present privacy leakage attacks that exploit metadata traces of web and research agents deployed locally. The attacks include (1) Prompt Recovery, which recovers a user’s original prompt from observed domain sequences, and (2) Trait Inference, which profiles latent attributes (e.g., gender, ideology) from the agent’s multi-session browsing patterns. The adversary trains an inference model using task-specific strategies such as In-Context Learning (ICL) and fine-tuning to map browsing traces to inference targets. Our results show that the attacks remain effective even when 40% of the traces are masked or fully obfuscated, revealing that encrypted content alone cannot guarantee privacy. We observe that WRAs visit dozens of marginal or redundant domains that do not improve the final report but still expand the adversary’s visibility. To mitigate these risks, we propose and evaluate practical defenses that either hide traces, by injecting LLM-generated decoy prompts and virtual personas to confuse inference, or block traces, by redirecting tasks to multipurpose sources or LLM knowledge, avoiding uniquely identifying domains.
@@
Benchmark dataset and OBELS metric: We construct and will release a dataset linking domain traces to prompts and persona traits, and introduce OBELS, a new ontology-aware metric for evaluating privacy leakage through behavioral traces.
Ablation and countermeasures: We conduct extensive experiments across agents, LLMs, and masking conditions, and evaluate mitigation strategies that constrain domain diversity or obfuscate traces while negligibly affecting utility.
@@ 2.1 AI, Web, and Research Agents
Standalone LLMs to AI Agents. LLMs excel at language understanding but are limited as static knowledge systems [2, 76]. They cannot access real-time information, call external APIs, or execute multi-step plans, but produce responses based only on their pretrained knowledge. Early solutions such as RAG [38, 24] and tool-augmented LLMs [63, 68, 60, 16] improved grounding and tool use, but remain constrained to pre-indexed corpora and invoking tools only when explicitly instructed. AI agents address this by embedding LLMs into a perception–action loop [86, 70, 79], where the model plans, executes, and adapts across multiple steps, enabling autonomous reasoning and dynamic interaction with external environments.
Web Agents. Generic agents can reason about which queries to run, but often lack robust capabilities for controlling browsers, parsing dynamic content, handling JavaScript-heavy pages, or managing session state across multiple sites. Web agents integrate LLM planning with browser control: issuing search queries, navigating between sources, extracting content, and deciding which actions to perform based on the evolving context [17]. Frameworks like AutoGen [84] and Browser-Use [43] exemplify these capabilities, though evaluations often emphasize narrow navigation tasks.
Research Web Agents (WRAs). WRAs operationalize deep evidence synthesis through a search–retrieve–extract–synthesize loop across many sources (Figure 2). Given a user query, the agent generates sub-queries and issues them via APIs such as Tavily [75] to minimize token overhead and retrieve relevant documents efficiently. Retrieved content is parsed into evidence, distilled into notes, and evaluated to decide whether further exploration is needed. This process repeats until sufficient coverage is achieved, at which point the agent composes a comprehensive, citation-rich output [30, 64, 90, 58, 17]. Systems exemplifying this paradigm include GPT Researcher [21], LangChain’s Local Deep Researcher [35], OpenAI Deep Research [55], Grok DeepSearch [85], and Gemini Deep Research [23].
Because WRAs generate dense browsing traces over dozens of domains, they amplify metadata-leakage risks. This risk has been underscored when Grok and OpenAI misconfigurations exposed private prompts to public indexing [40, 72]. Locally deployed agents can be one of the mitigation strategies against such threats. Running on personal machines or servers, they preserve confidentiality while offering lower latency, offline use, and lower cost. Examples include GPT-Researcher [21] with self-hosted backends such as Ollama [53] and AMD’s GAIA [3] to operate fully on-device, giving users more control over both their data and their agent workflows.
@@ 2.2 Privacy Attacks
Threats in LLMs. LLMs introduce unique privacy risks due to their training data and the ways users interact with them. One major class of attacks focuses on hidden-input inference, recovering sensitive inputs such as training data, system prompts, or user queries [12, 89, 31, 57, 39, 11, 82]. Other work has focused on recovering hidden system prompts used internally by providers [89, 31], where adversaries probe model input and output patterns to elicit protected instructions. Unlike these approaches, our setting assumes no direct access to the model and relies instead on indirect behavioral signals.
@@
Beyond direct extraction, studies show that semantic leakage occurs even when the attacker only observes generated outputs. Pasquini et al.[57] and Liu et al.[39] found that reconstructed prompts often preserve enough semantic content for intent inference, while Carlini et al.[11] and Weiss et al.[82] demonstrated that response timing and token-length patterns can reveal coarse-grained information about prompt topics. These attacks, however, assume access to model outputs or fine-grained response timings. By contrast, we show that passively observable network-level metadata alone, without accessing queries or responses, can be sufficient to infer hidden prompts.
@@
LLMs also introduce privacy risks through trait inference and user profiling. Their outputs carry author-identifying signals [45, 67, 28, 6], allowing inference of attributes such as age, gender, and personality. Other risks include speculative decoding revealing fragments of private inputs [81] and long-context memory retaining user data across sessions for later retrieval [78]. Unlike these works, we assume a strictly weaker threat model: the adversary does not observe model inputs, outputs, or internal states, relying solely on external traffic traces (e.g., domain sequences, timing, payload sizes).
@@
Threats in Web Privacy. Metadata leakage has long been recognized as a central privacy risk on the web. Classical work on anonymous communications (e.g., MIX Nets [15] and Tor [19]) demonstrated that encrypting content is insufficient when traffic patterns remain exposed. Website-fingerprinting studies [83, 9, 25, 33] showed that adversaries can infer visited pages based on encrypted flow patterns, while large-scale deanonymization attacks like the Netflix attack[46] revealed that sparse behavioral traces are often enough to re-identify individuals. Subsequent work confirmed that traffic features, such as packet sizes, burst patterns, and inter-arrival timing, can fingerprint websites even under protections like Tor [26, 56]. Parallel work in web tracking exposed persistent profiling vectors, including cookie-based tracking [37], cookieless fingerprinting [47, 1], and evercookies [4].
@@
Recent work has shifted to domain-level metadata as a source of behavioral leakage. Oliveira et al.[52] showed that as few as four domains can uniquely re-identify most users in real-world datasets. Crichton et al.[18] found that domain sequence fingerprints remain stable across years, even after cookie resets. Cross-context linkage is also possible: Naini et al.[44] and Su et al.[73] demonstrated that anonymized browsing logs can be matched across platforms, breaking assumed protections. Unlike fingerprinting, we attempt to reconstruct prompts or user traits that WRAs leak.
@@
While protocol-level defenses such as HTTPS, DNS-over-HTTPS (DoH), and DNS-over-TLS (DoT) aim to hide domain names too, they often fail to conceal key identifiers. Hoang et al.[27] showed that TLS handshakes frequently leak domain information, while Sunahara et al.[74] pointed out widespread misconfigurations in encrypted DNS deployments, enabling persistent metadata leakage. Similarly, Wang et al. [80] demonstrated that in-app telemetry (e.g., navigation events, click logs, and tracking APIs) leaks fine-grained behavioral patterns to platforms. Unlike these works, we assume a weaker adversary: a passive network observer without privileged telemetry access, relying only on network-level domain sequences and packet metadata.
@@
Agentic Privacy and Security Threats. AI agents amplify privacy risks by combining LLM reasoning with active browsing and external tool integration [20]. Kim et al. [34] showed that compromised agents can scrape PII from visited pages and reuse it in phishing, while InjecAgent [88] demonstrated that crafted payloads in webpages can trigger secret retrieval or unauthorized service access. Defenses such as AirGapAgent [7] and CHeaT [5] aim to counter these threats through context isolation, behavior cloaking, and honey-token traps.
@@ 3 Overview: Prompt and Trait Leakage Attack
Prompt Recovery. With proxy dataset consisting of pairs indexed by , the adversary constructs an inference model (via ICL or fine-tuning). For a new prompt with trace (optionally filtered to ), the adversary recovers the prompt , aiming to maximize functional similarity:
Trait Inference. Let denote a user’s latent trait vector and the prompts across sessions, yielding . From proxy dataset , the adversary instantiates an inference model (ICL). On a new multi-session trace and infers trait , maximizing agreement:

### v3

@@ Abstract
We show that Web and Research Agents (WRAs)—language-model-based systems that investigate complex topics on the Internet—are vulnerable to inference attacks by passive network observers. Deployment of WRAs locally by organizations and individuals for privacy, legal, or financial purposes exposes them to DNS resolvers, malicious ISPs, VPNs, web proxies, and corporate or government firewalls. However, unlike sporadic and scarce web browsing by humans, WRAs visit domains per each request with a distinct timing pattern creating unique privacy risks.
Specifically, we demonstrate a novel prompt and user trait leakage attack against WRAs that only leverages their network-level metadata (i.e., visited IP addresses and their timings). We start by building a new dataset of WRA traces based on real user search queries and queries generated by synthetic personas. We define a behavioral metric (called OBELS) to comprehensively assess similarity between original and inferred prompts, showing that our attack recovers over 73% of the functional and domain knowledge of user prompts. Extending to a multi-session setting, we recover up to 19 of 32 latent traits with high accuracy. Our attack remains effective under partial observability and noisy conditions. Finally, we discuss mitigation strategies that constrain domain diversity or obfuscate traces, showing negligible utility impact while reducing attack effectiveness by an average of 29%.
@@ 1 Introduction
Web and Research Agents (WRAs) are reshaping how users interact with information. Unlike traditional assistants that answer isolated queries, these Large Language Model (LLM)-powered systems autonomously plan, browse, and synthesize knowledge across the web [huang2025deep]. Major AI providers are integrating such agents directly into their flagship assistants or browsers, including OpenAI [openai2025introducingdeepresearch], Google Gemini [google2025geminideepresearch], Grok [xai2025grok3], Mistral [mistral2025lechat], and Perplexity [perplexity2025deepresearch]. Open-source counterparts such as LangChain’s Local Deep Researcher [langchain2025localdeepresearcher], GPT Researcher [elovic2023gptresearcher], and Agent Laboratory [schmidgall2025agent] are also widely adopted in academic and developer communities. Their growing deployment in assistants, productivity platforms, and enterprise environments, particularly within privacy-sensitive sectors such as healthcare, law, and finance, raises urgent concerns about their security and privacy implications. While on-device execution through locally-deployable models and TLS-encrypted access could suggest privacy guarantees, this perception is misleading. Yet WRAs leave unavoidable, structured metadata traces that can reveal sensitive user intent, creating an underexplored leakage vector.
As autonomous agent systems, WRAs inherit vulnerabilities from both LLMs and traditional web browsing [hui2024pleak, carlini2021extracting, siby2019encrypted, bird2020replication]. Yet they also expose a distinct, underexplored vulnerability: behavioral traces determined by an LLM acting on a user’s query. Prior work has shown that user intent and goals can be inferred from search queries and browsing activity [rose2004understanding, jones2007know]. Even when content is encrypted, metadata such as domain names, access order, payload size, and timing can silently reveal what a user is trying to accomplish [oliveira2023browsing, su2017anonymizing].
Crucially, these behavioral traces are operationally distinguishable. WRAs generate dense cascades of semantically related domain visits within short time windows [perplexity2025deepresearch]. Unlike cognitively paced, irregular, and opportunistic human browsing [kellar2007field], WRAs follow deterministic retrieval-and-synthesis loops with low variance in per-step timing. As a result, a single WRA session often traverses 40–100 distinct domains, whereas human sessions typically rely on few sites [wang2025ai]. These structured, high-throughput access patterns make WRA traffic reliably separable from human activity and particularly amenable to profiling and inference attacks [chiang2025web, mudryi2025hidden].
@@
We adopt a realistic threat model in which a passive adversary, (e.g., DNS resolver, ISP, enterprise firewall, or local network operator) observes only domain-level traffic metadata, without access to prompts, page contents, or model internals. Despite this limited visibility, we show that adversaries can reliably recover what users asked and who they are.
@@
Motivating Examples. Consider a user seeking reproductive health information: their research agent may visit clinics, support forums, and medical sites. Even without access to page contents, these traces alone can expose the agent’s internal reasoning process and deeply stigmatized intent. As WRAs become integrated into daily decision-making, their behaviors present a new leakage channel. In essence, privacy failures emerge not from what agents say, but from how they act.
@@
Our Work. We present privacy leakage attacks that exploit metadata traces of WRAs deployed locally. The attacks include (1) Prompt Recovery, which recovers a user’s original prompt from observed domain sequences, and (2) Trait Inference, which profiles latent attributes (e.g., gender, ideology) from the agent’s multi-session browsing patterns. The adversary trains an inference model using task-specific strategies such as In-Context Learning (ICL) and fine-tuning to map browsing traces to inference targets (Table 2). Our results show that the attacks remain effective even when 40% of the traces are masked or fully obfuscated, revealing that encrypted content alone cannot guarantee privacy (Table 10). We observe that WRAs visit dozens of marginal or redundant domains that do not improve the final report but still expand the adversary’s visibility (Table 11). To mitigate these risks, we propose and evaluate practical defenses that either hide traces, by injecting LLM-generated decoy prompts and virtual personas to confuse inference (Section 8.1), or block traces, by redirecting tasks to multipurpose sources or LLM knowledge, avoiding uniquely identifying domains (Section 8.2).
@@
Benchmark dataset and OBELS metric: We construct and release a dataset linking domain traces to prompts and persona traits, and introduce OBELS, an ontology-aware metric for evaluating privacy leakage through behavioral traces.
Ablation and mitigation: We conduct extensive experiments across agents, LLMs, and masking conditions, and evaluate mitigation that constrain domain diversity or obfuscate traces while negligibly affecting utility.
@@ 2.1 AI, Web, and Research Agents
Standalone LLMs to AI Agents. LLMs excel at language understanding but are limited as static knowledge systems [achiam2023gpt, team2024gemini]. They cannot access real-time information, call external APIs, or execute multi-step plans, but produce responses based only on their pretrained knowledge. Early solutions such as RAG [lewis2020retrieval, guu2020retrieval] and tool-augmented LLMs [schick2023toolformer, shen2023hugginggpt, qin2023toolllm, chen2024advancing] improved grounding and tool use, but remain constrained to pre-indexed corpora and invoking tools only when explicitly instructed. AI agents address this by embedding LLMs into a perception–action loop [yao2023react, autogpt, wang2023voyager], where the model plans, executes, and adapts across multiple steps, enabling autonomous reasoning and dynamic interaction with external environments.
Web Agents. Generic agents can reason about which queries to run, but often lack robust capabilities for controlling browsers, parsing dynamic content, handling JavaScript-heavy pages, or managing session state across multiple sites. Web agents integrate LLM planning with browser control: issuing search queries, navigating between sources, extracting content, and deciding which actions to perform based on the evolving context [chiang2025web]. Frameworks like AutoGen [wu2023autogen] and Browser-Use [browser_use2024] exemplify these capabilities, though evaluations often emphasize narrow navigation tasks.
Research Web Agents (WRAs). WRAs operationalize deep evidence synthesis through a search–retrieve–extract–synthesize loop across many sources (Figure 2). Given a user query, the agent generates sub-queries and issues them via APIs such as Tavily [tavily] to minimize token overhead and retrieve relevant documents efficiently. Retrieved content is parsed into evidence, distilled into notes, and evaluated to decide whether further exploration is needed. This process repeats until sufficient coverage is achieved, at which point the agent composes a comprehensive, citation-rich output [huang2025deep, schmidgall2503agentrxiv, zheng2025deepresearcher, perplexity2025deepresearch, chiang2025web]. Systems exemplifying this paradigm include GPT Researcher [elovic2023gptresearcher], LangChain’s Local Deep Researcher [langchain2025localdeepresearcher], OpenAI Deep Research [openai2025introducingdeepresearch], Grok DeepSearch [xai2025grok3], and Gemini Deep Research [google2025geminideepresearch].
Because WRAs generate dense browsing traces over dozens of domains, they amplify metadata-leakage risks. This risk has been underscored when Grok and OpenAI misconfigurations exposed private prompts to public indexing [bbc2025example, stokel-walker2025google]. Locally deployed agents can be one of the mitigation strategies against such threats. Running on personal machines or servers, they preserve confidentiality while offering lower latency, offline use, and lower cost. Examples include GPT-Researcher [elovic2023gptresearcher] with self-hosted backends such as Ollama [ollama2025] and AMD’s GAIA [amd_gaia_github] to operate fully on-device, giving users more control over both their data and their agent workflows.
@@ 2.2 Privacy Attacks
Threats in LLMs. LLMs introduce unique privacy risks due to their training data and the ways users interact with them. One major class of attacks focuses on hidden-input inference, recovering sensitive inputs such as training data, system prompts, or user queries [carlini2021extracting, zhang2023effective, hui2024pleak, pasquini2024llmmap, liu2025evaluatingllm, carlini2024remote, weiss2024your]. Other work has focused on recovering hidden system prompts used internally by providers [zhang2023effective, hui2024pleak], where adversaries probe model input and output patterns to elicit protected instructions. Unlike these approaches, our setting assumes no direct access to the model and relies instead on indirect behavioral signals.
@@
Beyond direct extraction, studies show that semantic leakage occurs even when the attacker only observes generated outputs. Pasquini et al.[pasquini2024llmmap] and Liu et al.[liu2025evaluatingllm] found that reconstructed prompts often preserve enough semantic content for intent inference, while Carlini et al.[carlini2024remote] and Weiss et al.[weiss2024your] demonstrated that response timing and token-length patterns can reveal coarse-grained information about prompt topics. These attacks, however, assume access to model outputs or fine-grained response timings. By contrast, we show that passively observable network-level metadata alone, without accessing queries or responses, can be sufficient to infer hidden prompts.
@@
LLMs also introduce privacy risks through trait inference and user profiling. Their outputs carry author-identifying signals [narayanan2012feasibility, schwartz2013personality, hovy2015demographic, azucar2018predicting], allowing inference of attributes such as age, gender, and personality. Other risks include speculative decoding revealing fragments of private inputs [wei2024privacy] and long-context memory retaining user data across sessions for later retrieval [wang2025unveiling]. Unlike these works, we assume a strictly weaker threat model: the adversary does not observe model inputs, outputs, or internal states, relying solely on external traffic traces (e.g., domain sequences, timing, payload sizes).
@@
Threats in Web Privacy. Metadata leakage has long been recognized as a central privacy risk on the web. Classical work on anonymous communications (e.g., MIX Nets [chaum1981untraceable] and Tor [dingledine2004tor]) demonstrated that encrypting content is insufficient when traffic patterns remain exposed. Website Fingerprinting (WF) studies [wright2009traffic, cai2012touching, hayes2016k, juarez2016toward] showed that adversaries can infer visited pages based on encrypted flow patterns, while large-scale deanonymization attacks like the Netflix attack[narayanan2008robust] revealed that sparse behavioral traces are often enough to re-identify individuals. Subsequent work confirmed that traffic features, such as packet sizes, burst patterns, and inter-arrival timing, can fingerprint websites even under protections like Tor [hintz2002fingerprinting, panchenko2011website]. Parallel work in web tracking exposed persistent profiling vectors, including cookie-based tracking [lerner2016internet], cookieless fingerprinting [nikiforakis2013cookieless, acar2014web], and evercookies [ayenson2011flash].
@@
Recent work has shifted to domain-level metadata as a source of behavioral leakage. Oliveira et al.[oliveira2023browsing] showed that as few as four domains can uniquely re-identify most users in real-world datasets. Crichton et al.[crichton2025rethinking] found that domain sequence fingerprints remain stable across years, even after cookie resets. Cross-context linkage is also possible: Naini et al.[naini2015you] and Su et al.[su2017anonymizing] demonstrated that anonymized browsing logs can be matched across platforms, breaking assumed protections. Whereas classical WF identifies which site a user visits, our attack infers user’s task or intent from agent-driven traversals. The two are complementary: WF reveals destinations, while our method reconstructs prompts or traits from the resulting domain sequences.
@@
While protocol-level defenses such as HTTPS, DNS-over-HTTPS (DoH), and DNS-over-TLS (DoT) aim to hide domain names too, they often fail to conceal key identifiers. Hoang et al.[hoang2020assessing] showed that TLS handshakes frequently leak domain information, while Sunahara et al.[sunahara2025framework] pointed out widespread misconfigurations in encrypted DNS deployments, enabling persistent metadata leakage. Similarly, Wang et al. [wang2025wechat] demonstrated that in-app telemetry (e.g., navigation events, click logs, and tracking APIs) leaks fine-grained behavioral patterns to platforms. Unlike these works, we assume a weaker adversary: a passive network observer without privileged telemetry access, relying only on network-level domain sequences and packet metadata.
@@
Agentic Privacy and Security Threats. AI agents amplify privacy risks by combining LLM reasoning with active browsing and external tool integration [googlereport]. Kim et al. [kim2024llms] showed that compromised agents can scrape PII from visited pages and reuse it in phishing, while InjecAgent [zhan2024injecagent] demonstrated that crafted payloads in webpages can trigger secret retrieval or unauthorized service access. Defenses such as AirGapAgent [bagdasarian2024airgapagent] and CHeaT [ayzenshteyncloak] aim to counter these threats through context isolation, behavior cloaking, and honey-token traps.
@@ 3 Overview: Prompt and Trait Leakage Attack
Prompt Recovery. With proxy dataset consisting of pairs indexed by , the adversary constructs an inference model . For a new prompt with trace (optionally filtered to ), the adversary recovers the prompt , aiming to maximize functional similarity:
Trait Inference. Let denote a user’s latent trait vector and the prompts across sessions, yielding . From proxy dataset , the adversary instantiates an inference model . On a new multi-session trace and infers trait , maximizing agreement: